A furniture manufacturer predicts that they will sell 12,000 of product A and 8,000 of product B in the next financial period. They prepare their budget accordingly.

At the end of the financial period the actual figures are 15,000 for product A and 7,000 for product B. Costs are assigned and the wholesale margin on product A is calculated to be $450 and on product B it is $350.

Calculate the predicted and the actual sales mix, the variances that need to be examined and their impact
1
Expert's answer
2018-08-23T11:14:08-0400
Predicted: 12,000 of product A and 8,000 of product B.
Actual: 15,000 for product A and 7,000 for product B.
The wholesale margin on product A is calculated to be $450 and on product B it is $350.
The predicted sales mix is 12,000/20,000 = 60% of A to 40% of B.
The actual sales mix is 15,000/22,000 = 68.2% of A to 31.8% of B.
The sales mix variance for A is 15,000*(0.682% - 0.6%)*$450 = 553,500 or favorable variance.
For B, the sales mix variance is 7,000*(0.318 - 0.4)*350 = -200,900 or unfavorable variance.